OpenStudy (firejay5):

Simplify. Show Work!!!!!! 18. 5t - 5 over t^2 - 1 19. c + 5 over 2c + 10

OpenStudy (raden):

hint : 5t-5 = 5(t-1) t^2 - 1 = (t+1)(t-1)

OpenStudy (raden):

cancel out t+1

OpenStudy (raden):

t-1, i meant :)

OpenStudy (firejay5):

The Answer is 5 over t + 1

OpenStudy (raden):

yup

OpenStudy (firejay5):

I just need work to show it

OpenStudy (raden):

5t - 5/ t^2 - 1 = 5(t-1) / (t+1)(t-1) = ur answer above

OpenStudy (firejay5):

What about 19

OpenStudy (raden):

do same idea for Q)19 factor out 2c+10 be 2(c+5) cancel out the c+5, get ur answer

OpenStudy (firejay5):

I have a question

OpenStudy (raden):

what is it ?

OpenStudy (firejay5):

Whenever I simplify, do I have to show my work, unless the teacher tells you to

OpenStudy (raden):

actually, YES! show and explain it so that ur techer happy :)

OpenStudy (firejay5):

I got 1/2 for question 19

OpenStudy (raden):

exactly, that's right

OpenStudy (firejay5):

I got my answer, it's just sometimes it's hard for me to show my work

OpenStudy (firejay5):

on question 19 how is it 1/2, is it because when you cancel c + 5 then the numerator is 1 and the 2 is on the denominator

OpenStudy (raden):

exercises by continue...